//Apiary Record 7th July 2019

Apiary Record 7th July 2019

Sixteen members and one guest attended the Apiary on muggy overcast afternoon, temperature 20C.

Last week queen cells were removed from four of the colonies, this week three of those colonies had swarmed.

Hive 1 – Anne C Ballard, Liz Hawes – WBC Double Brood. Swarmed, one good QC, position marked, 5/6 damaged QC’s.  B1: 3 brood, 2 stores, B2: 5 brood 2 stores, sufficient space & stores.  3 supers, 1 full 2 empty.  Some wax moth.  No inspection for 3 weeks

Hive 1B – David Parkinson – National 14X12, 5BIAS, 3 stores, sufficient space & stores.  Temper 8

Hive 1C – Not inspected, hatching queen. Next inspection >21 July. 

Hive 2 – Rowan Roberts, Ian & Cordi Hobbs – National Brood & 1/2.  Swarmed.  1 sealed, 1 open QC left in 1/2 brood, (other QC’s used to split colony?).  B1: 8 brood, B2: 4 brood. 10 stores. Sufficient room & stores.  2 Supers.  temper 8.  No inspection for 3 weeks

Hive 3 – Peter reader – National deep.  Swarmed.  1Sealed, 1 open QC kept, 5 removed.  Introduced 3 frames of brood from 3 other hives to boost depleted colony.  6 brood, 2 stores, sufficient room & stores.  2 supers.  Temper 7.  No inspection for 3 weeks

Hive 4 – Alex Nolan – National Deep, Queen seen (red), 6BIAS, 4 stores, 1 frame removed to reinforce nuc (?).  Sufficient room & stores.  4 supers.  Temper 8

Hive 5 – Richard Williams – National Double Brood undergoing a Bailey Comb Exchange.  Old Deep removed since all worker brood hatched & bees storing pollen in vacated cells.  Queen seen.  9 BIAS, 2 stores, short of space so 3 frames BIAS moved up into new Deep added to brood to increase space.  2 supers, 1 added.  Temper 9

Hive Q98 – Chris Clark, Gareth Cornelius – National Deep, queen seen, building well, 7 BIAS, one donated to Hive 3, 5 stores.  2 empty frames moved inside pollen barrier.  Sufficient room & stores.  1 super.  Temper 8

Hive Q96 – Chris Clark, Gareth Cornelius – National Deep, Queen seen, building well, 8 BIAS, one donated to Hive 3, 5 stores, 2 empty frames moved inside pollen barrier.  Sufficient room & stores. 1 super.  Temper 8

Gareth, Green Hive – Chris Clark, Gareth Cornelius – National 14X12, Queen seen, 5 BIAS, sufficient room & stores.  Super added.  New feeder added with eke.  Temper 10, very calm

Paul & Rachel Hive – not inspected, Virgin Queen

Horton Hive – Anne C Ballard, Steve Wilson – National Deep, Queen seen, 3 QC’s .  To prevent swarming Queen on brood frame + 1 stores moved to nuc on top of hive.  Two empty (?) QC’s removed, 1 open QC position marked, retained.  8 BIAS, 1 stores.  1 super.  Next inspection 7 days, remove all but 1 open QC & retain marked capped QC

By |2019-08-14T21:42:07+00:00July 10th, 2019|Apiary|0 Comments

Leave A Comment